What is the derivative of a + 15 over the square root of a?
Answer by stanbon(75887)

What is the derivative of a + 15 over the square root of a?
----------------
f(a) = (a+15)/a
---
Treating it as a quotient you get:
f'(a) = [a(1)-(a+15)(1)]/a^2 = -15/a^2
----
Changing the form you could write:
f(a) = 1+(15/a) = 1+15*a^-1
Then f'(a) = -15*a^-2 = -15/a^2
-------------
Cheers,
Stan H.